OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the equation 15x - 5y = 30. x-intercept = (-2, 0); y-intercept = (0, 6) x-intercept = (2, 0); y-intercept = (0, -6) x-intercept = (0, -6); y-intercept = (2, 0) x-intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (anonymous):

You want the answer, or how to get it?

OpenStudy (anonymous):

answer , please (:

OpenStudy (anonymous):

Ok :) the answer is x-intercept = (2, 0); y-intercept = (0, -6)

OpenStudy (anonymous):

THANK YOU SO MUCH <333 :D

OpenStudy (anonymous):

15x - 5y = 30 15(0) - 5y = 30 -5y = 30 -5y/-5 = 30/-5 y = -6 15x - 5(0) = 30 15x = 30 15x/15 = 30/15 x = 2 (2,-6) no prob

OpenStudy (anonymous):

Thanks for the medal